OpenStudy (marysammonss):

A line has slope 5/6 and y–intercept −3−3. Which answer is the equation of the line? y=3x+5/6 y=5/6x+3y y=−3x+5/6 y=5/6x−3

OpenStudy (marysammonss):

A line has slope 5/6 and y–intercept −3 Which answer is the equation of the line? y=3x+5/6 y=5/6x+3y y=−3x+5/6 y=5/6x−3